\begin{thebibliography}{99}

\bibitem{ReWh} T. Regge and J. A. Wheeler, Phys. Rev. {\bf 108}, 1063 (1957).

\bibitem{Nollert1} H. P. Nollert, Class. Quantum Grav. {\bf 16}, R159 (1999); E. Berti, V. Cardoso and A. O. Starinets, Class. Quant. Grav. {\bf
26}, 163001 (2009).

\bibitem{Press} W. H. Press, Astrophys. J. {\bf 170}, L105 (1971);
V. de la Cruz, J. E. Chase and W. Israel, Phys. Rev. Lett. {\bf 24},
423 (1970); C.V. Vishveshwara, Nature {\bf 227}, 936 (1970); M.
Davis, R. Ruffini, W. H. Press and R. H. Price, Phys. Rev. Lett.
{\bf 27}, 1466 (1971).

\bibitem{Tails} R.H. Price, Phys. Rev. D {\bf 5}, 2419 (1972);
C. Gundlach, R.H. Price, and J. Pullin, Phys. Rev. D {\bf 49}, 883
(1994); J. Bic\'ak, Gen. Relativ. Gravitation {\bf 3}, 331 (1972);
E. S. C. Ching, P. T. Leung, W. M. Suen, and K. Young, Phys. Rev.
Lett. {\bf 74}, 2414 (1995); E. S. C. Ching, P. T. Leung, W. M.
Suen, and K. Young, Phys. Rev. D {\bf 52}, 2118 (1995); S. Hod and
T. Piran, Phys. Rev. D {\bf 58}, 024017 (1998)
[arXiv:gr-qc/9712041]; S. Hod and T. Piran, Phys. Rev. D {\bf 58},
024018 (1998) [arXiv:gr-qc/9801001]; S. Hod and T. Piran, Phys. Rev.
D {\bf 58}, 044018 (1998) [arXiv:gr-qc/9801059]; S. Hod and T.
Piran, Phys. Rev. D {\bf 58}, 024019 (1998) [arXiv:gr-qc/9801060];
S. Hod, Phys. Rev. D {\bf 58}, 104022 (1998) [arXiv:gr-qc/9811032];
S. Hod, Phys. Rev. D {\bf 61}, 024033 (2000) [arXiv:gr-qc/9902072];
S. Hod, Phys. Rev. D {\bf 61}, 064018 (2000) [arXiv:gr-qc/9902073];
L. Barack, Phys. Rev. D {\bf 61}, 024026 (2000); S. Hod, Phys. Rev.
Lett. {\bf 84}, 10 (2000) [arXiv:gr-qc/9907096]; S. Hod, Phys. Rev.
D {\bf 60}, 104053 (1999) [arXiv:gr-qc/9907044]; S. Hod, Class.
Quant. Grav. {\bf 26}, 028001 (2009) [arXiv:0902.0237]; S. Hod,
Class. Quant. Grav. {\bf 18}, 1311 (2001) [arXiv:gr-qc/0008001]; S.
Hod, Phys. Rev. D {\bf 66}, 024001 (2002) [arXiv:gr-qc/0201017]; R.
J. Gleiser, R. H. Price, and J. Pullin, Class. Quant. Grav. {\bf
25}, 072001 (2008); M. Tiglio, L. E. Kidder, and S. A. Teukolsky,
Class. Quant. Grav. {\bf 25}, 105022 (2008); R. Moderski and M.
Rogatko, Phys. Rev. D {\bf 77}, 124007 (2008); X. He and J. Jing,
Nucl. Phys. B {\bf 755}, 313 (2006); H. Koyama and A. Tomimatsu,
Phys. Rev. D {\bf 65}, 084031 (2002); B. Wang, C. Molina, and E.
Abdalla, Phys. Rev. D {\bf 63}, 084001 (2001); S. Hod, Phys. Rev.
Lett. {\bf 81}, 4293 (1998) [arXiv:gr-qc/9812002]; G. T. Horowitz
and V. E. Hubeny, Phys. Rev. D {\bf 62}, 024027 (2000); E. W.
Leaver, Proc. R. Soc. A {\bf 402}, 285 (1985); S. Hod, Phys. Rev. D
{\bf 75}, 064013 (2007) [arXiv:gr-qc/0611004]; S. Hod, Class. and
Quant. Grav. {\bf 24}, 4235 (2007) [arXiv:0705.2306]; A. Gruzinov,
arXiv:gr-qc/0705.1725; A. Pesci, Class. Quantum Grav. {\bf 24}, 6219
(2007); S. Hod, Phys. Rev. D {\bf 78}, 084035 (2008); S. Hod, Phys.
Lett. B {\bf 666} 483 (2008) [arXiv:0810.5419]; S. Hod, Phys. Rev. D
{\bf 80}, 064004 (2009) [arXiv:0909.0314]; S. Hod, Phys. Lett. A
{\bf 374}, 2901 (2010) [arXiv:1006.4439]; S. Hod, Phys. Rev. D. {\bf
84}, 044046 (2011) [arXiv:1109.4080]; S. Hod, Phys. Lett. B {\bf
710}, 349 (2012) [arXiv:1205.5087]; B. Mashhoon, Phys. Rev. D {\bf
31}, 290 (1985); C. J. Goebel, Astrophys. J. {\bf 172}, L95 (1972);
V. Cardoso, A. S. Miranda, E. Berti, H. Witek, and V. T. Zanchin,
Phys. Rev. D {\bf 79}, 064016 (2009); S. R. Dolan and A. C.
Ottewill, Classical Quantum Gravity {\bf 26}, 225003 (2009); S. R.
Dolan, Phys. Rev. D {\bf 82}, 104003 (2010); Y. D\'ecanini, A.
Folacci, and B. Raffaelli, Phys. Rev. D {\bf 84}, 084035 (2011).

\bibitem{Chan} S. Chandrasekhar, {\it The Mathematical Theory of Black
Holes}, (Oxford University Press, New York, 1983).

\bibitem{Kerr} R. P. Kerr, Phys. Rev. Lett. {\bf 11}, 237 (1963).

\bibitem{Teu} S. A. Teukolsky, Phys. Rev. Lett. {\bf 29}, 1114 (1972);
Astrophys. J. {\bf 185}, 635 (1973).

\bibitem{Heun} A. Ronveaux, {\it Heun's differential equations}.
(Oxford University Press, Oxford, UK, 1995).

\bibitem{Fiz1} P. P. Fiziev, e-print arXiv:0902.1277; R. S. Borissov and P. P. Fiziev, e-print arXiv:0903.3617;
P. P. Fiziev, Phys. Rev. D {\bf 80}, 124001 (2009); P. P. Fiziev,
Class. Quant. Grav. {\bf 27}, 135001 (2010).

\bibitem{Abram} M. Abramowitz and I. A. Stegun, {\it Handbook of
Mathematical Functions} (Dover Publications, New York, 1970).

\bibitem{Flam} C. Flammer, {\it Spheroidal Wave Functions} (Stanford
University Press, Stanford, 1957).

\bibitem{Meix} J. Meixner and F. W. Sch\"afke, {\it Mathieusche Funktionen und Sph\"aroidfunktionen
mit Anwendungen auf Physikalische und Technische Probleme}
(Springer-Verlag, Berlin, 1954).

\bibitem{Breu1} R. A. Breuer, {\it Gravitational Perturbation Theory and Synchrotron Radiation}
(Lecture Notes in Physics, Vol. 44), (Springer, Berlin 1975).

\bibitem{Breu2} R. A. Breuer, M. P. Ryan Jr, and S. Waller, Proc. R. Soc. London {\bf A358}, 71 (1977).

\bibitem{Cas} M. Casals and A. C. Ottewill, Phys. Rev. D {\bf 71}, 064025 (2005).

\bibitem{BCC} E. Berti, V. Cardoso, and M. Casals, Phys. Rev. D {\bf 73}, 024013 (2006);
Erratum-ibid. D {\bf 73}, 109902 (2006).

\bibitem{Hodo} S. Hod, Phys. Lett. B {\bf 717}, 462 (2012) [arXiv:1304.0529].

\bibitem{HodL1} S. Hod, Phys. Rev. Lett. {\bf 81}, 4293 (1998) [arXiv:gr-qc/9812002].

\bibitem{HodL2} U. Keshet and S. Hod, Phys. Rev. D {\bf 76}, R061501 (2007) [arXiv:0705.1179].

\bibitem{BerCarY} E. Berti, V. Cardoso and S. Yoshida, Phys. Rev. D {\bf 69}, 124018 (2004).

\bibitem{Yang} H. Yang, D. A. Nichols, F. Zhang, A. Zimmerman, Z. Zhang and Y.
Chen, arXiv:1207.4253.

\bibitem{NoteSch} Note that in the quantum-mechanical terminology
$-U$ stands for ${{2m}\over{\hbar^2}}(E-V)$, where $E, V$, and $m$
are the total energy, potential energy, and mass of the of the
particle, respectively.

\bibitem{WKB1} L. D. Landau and E. M. Liftshitz, {\it Quantum Mechanics}, 3rd ed. (Pergamon, New York,
1977), Chap. VII.

\bibitem{WKB2} J. Heading, {\it An Introduction to Phase Integral Methods} (Wiley,
New York, 1962).

\bibitem{WKB3} C. M. Bender and S. A. Orszag, {\it Advanced Mathematical
Methods for Scientists and Engineers} (McGraw-Hill, New York, 1978),
Chap. 10.

\bibitem{Iyer} B. F. Schutz and C. M. Will, Astrophys. J. {\bf 291}, L33 (1985);
S. Iyer and C. M. Will, Phys. Rev. D {\bf 35}, 3621 (1987).

\bibitem{Notehigh} Higher order corrections to the asymptotic eigenvalues [see formula (\ref{Eq18}) below] can
be obtained by using a higher-order WKB analysis \cite{Iyer}.

%\bibitem{Zhou} C. S. Park, M. G. Jeong, S. K. Yoo, and D. K. Park, Phys.
%Rev. A, {\bf 58}, 3443 (1998); Z. Cao, Q. Liu, Q. Shen, X. Dou, Y.
%Chen, and Y. Ozaki, Phys. Rev. A {\bf 63}, 054103 (2001); F. Zhou,
%Z. Cao, and Q. Shen, Phys. Rev. A {\bf 67}, 062112 (2003).

\bibitem{Noteapp} Here we have used the relations
$\sin\theta=1-{1\over 2}(\theta-{\pi\over 2})^2+O[(\theta-{\pi\over
2})^4]$ and $\cos\theta=-(\theta-{\pi\over 2})+O[(\theta-{\pi\over
2})^3]$.
%We shall restrict ourselves to the leading-order asymptotic
%behavior of the eigenvalues in the $c_I\to\infty$ limit and thus we
%may neglect terms of order $O(1)$ and smaller such as

\bibitem{NoteCas} Note that our discrete resonance parameter $N$ corresponds
to $l-\text{max}(|m|,|s|)$ of \cite{Flam,Meix}.

\bibitem {Miller} P. D. Miller, Physica D {\bf 152–153}, 145 (2001).

\end{thebibliography}

\documentclass[preprintnumbers,amsmath,amssymbm,prd]{revtex4}
\usepackage{epsfig}
\usepackage{graphicx}
\usepackage{amssymb}

\begin{document}
\title{Black-hole perturbation theory: The asymptotic spectrum of the prolate spin-weighted spheroidal
harmonics}
\author{Shahar Hod}
\address{The Ruppin Academic Center, Emeq Hefer 40250, Israel}
\address{ }
\address{The Hadassah Institute, Jerusalem 91010, Israel}
\date{\today}

\begin{abstract}
\ \ \ Prolate spin-weighted spheroidal harmonics play a key role in
black-hole perturbation theory. In particular, the highly damped
quasinormal resonances of rotating Kerr black holes are closely
related to the asymptotic eigenvalues of these important functions.
We here present a novel and compact derivation of the asymptotic
eigenvalues of the prolate spin-weighted spheroidal harmonics. Our
analysis is based on a simple trick which transforms the
corresponding spin-weighted spheroidal angular equation into a
Schr\"odinger-like wave equation which is amenable to a standard WKB
analysis. Our {\it analytical} results for the prolate asymptotic
spectrum agree with previous {\it numerical} computations of the
eigenvalues which appear in the literature.
\end{abstract}
\bigskip
\maketitle

%]

\section{Introduction.}

The characteristic dynamics of test fields in black-hole spacetimes
has been studied extensively since the pioneering work of Regge and
Wheeler \cite{ReWh}, see also \cite{Nollert1,Press,Tails} and
references therein. An astrophysically realistic model of wave
dynamics in black-hole spacetimes should involve a non-spherical
background geometry with angular momentum. In terms of the
Boyer-Lindquist coordinates, the spacetime of a rotating Kerr black
hole is described by the line-element \cite{Chan,Kerr}
\begin{eqnarray}\label{Eq1}
ds^2=-\Big(1-{{2Mr}\over{\rho^2}}\Big)dt^2-{{4Mar\sin^2\theta}\over{\rho^2}}dt
d\phi+{{\rho^2}\over{\Delta}}dr^2
%\nonumber \\
+\rho^2d\theta^2+\Big(r^2+a^2+{{2Ma^2r\sin^2\theta}\over{\rho^2}}\Big)\sin^2\theta
d\phi^2,
\end{eqnarray}
where $M$ and $a$ are the mass and angular momentum per unit mass of
the black hole, respectively. (We use gravitational units in which
$G=c=1$). Here $\Delta\equiv r^2-2Mr+a^2$ and $\rho\equiv
r^2+a^2\cos^2\theta$.

In this paper we consider perturbations of the {\it non}-spherical
Kerr spacetime. The dynamics of a test field $\Psi$ in the rotating
Kerr spacetime is governed by the well-known Teukolsky master
equation \cite{Teu}. One may decompose the field as
\begin{equation}\label{Eq2}
_s\Psi_{lm}(t,r,\theta,\phi)=e^{im\phi}{_sS_{lm}}(\theta;a\omega){_s\psi_{lm}}(r)e^{-i\omega
t}\ ,
\end{equation}
where $\omega$ is the (conserved) frequency of the mode, $l$ is the
spheroidal harmonic index, and $m$ is the azimuthal harmonic index.
The parameter $s$ is called the spin weight of the field, and is
given by $s=\pm 2$ for gravitational perturbations, $s=\pm 1$ for
electromagnetic perturbations, $s=\pm {1\over 2}$ for massless
neutrino perturbations, and $s=0$ for scalar perturbations
\cite{Teu}.

With the decomposition (\ref{Eq2}), $\psi$ and $S$ obey radial and
angular equations, both of confluent Heun type
\cite{Teu,Heun,Fiz1,Abram,Flam}, coupled by a separation constant
$A(a\omega)$. The radial Teukolsky equation is given by \cite{Teu}
\begin{equation}\label{Eq3}
\Delta^{-s}{{d}
\over{dr}}\Big(\Delta^{s+1}{{d\psi}\over{dr}}\Big)+\Big[{{K^2-2is(r-M)K}\over{\Delta}}
-a^2\omega^2+2ma\omega-A+4is\omega r\Big]\psi=0\ ,
\end{equation}
where $K\equiv (r^2+a^2)\omega-am$.

The angular functions $S(\theta;a\omega)$ are the spin-weighted
spheroidal harmonics which are solutions of the angular equation
\cite{Teu,Heun,Fiz1,Abram,Flam}
\begin{equation}\label{Eq4}
{1\over {\sin\theta}}{\partial \over
{\partial\theta}}\Big(\sin\theta {{\partial
S}\over{\partial\theta}}\Big)+\Big[c^2\cos^2\theta-2c
s\cos\theta-{{(m+s\cos\theta)^2}\over{\sin^2\theta}}+s+A\Big]S=0\  ,
\end{equation}
where $c\equiv a\omega$.
%\in\mathbb{C}$
The angular functions are
required to be regular at the poles $\theta=0$ and $\theta=\pi$.
These boundary conditions pick out a discrete set of eigenvalues
$\{_sA_{lm}\}$ labeled by the integers $m$ and $l$. [In the $c\to 0$
limit these angular functions become the familiar spin-weighted
spherical harmonics with the corresponding angular eigenvalues
$A=l(l+1)-s(s+1)+O(a\omega)$.]

The spin-weighted spheroidal harmonics $S(\theta;c)$ and their
corresponding eigenvalues $\{_sA_{lm}\}$ have attracted much
attention over the years from both physicists and mathematicians
\cite{Flam,Meix,Breu1,Breu2,Cas,BCC,Hodo}. It is worth emphasizing
that in order to compute the characteristic resonances of black
holes \cite{Nollert1}, one must {\it first} compute the (closely
related) angular eigenvalues $\{_sA_{lm}\}$ [see Eq. (\ref{Eq3})].
In particular, in the framework of semi-classical general
relativity, it has been conjectured that the highly damped
resonances may shed light on the quantum properties of black holes
\cite{HodL1,HodL2}. For rotating black holes, these highly damped
resonances are characterized by $c_I\to\infty$ (where $c_I\equiv \Im
c$).

The asymptotic limit $c_I\to\infty$ of the angular equation
(\ref{Eq4}) was studied in \cite{Flam,Meix} for the $s=0$ case. It
was found that the asymptotic scalar eigenvalues are given by
\begin{equation}\label{Eq5}
_0A_{lm}=[2(l-|m|)+1] |c_I|+O(1)\  .
\end{equation}
As pointed out in \cite{BCC}, the analysis of \cite{Flam,Meix} for
the spin-$0$ case is somewhat incomplete. The analysis of
\cite{Flam,Meix} requires the knowledge of the number of zeros of
the scalar harmonics in the interval $[0,\pi]$. However, as
emphasized in \cite{BCC}, the approximated solution found in
\cite{Flam,Meix} for the $c_I\to\infty$ limit is only valid in a
region far from the end-points $\theta=0,\pi$. Thus, the analysis of
\cite{Flam,Meix} cannot rule out the possible existence of
additional zeros of the angular eigenfunctions near the end-points.
The possible omission of such zeros would lead to a wrong asymptotic
behavior of the prolate (with $c=ic_I$) eigenvalues, see \cite{BCC}
for details. In this respect the analysis of \cite{Flam,Meix} is not
complete. One of the goals of the present paper is to present a more
rigorous proof of the formula (\ref{Eq5}) for the asymptotic scalar
eigenvalues.
%Moreover, we
%believe that our proof is also simpler and more intuitive than the
%one given in \cite{Flam,Meix} for the spin-$0$ case.

The asymptotic spectrum of the prolate eigenvalues for the general
spin case was first studied in \cite{Breu1}. However, as pointed out
in \cite{BCC,BerCarY}, the analysis of \cite{Breu1} for the general
spin case is fundamentally flawed (see \cite{BCC,BerCarY} for
details). The {\it numerical} results presented in Refs.
\cite{BCC,BerCarY} for the general spin case provide evidence for
the asymptotic behavior
\begin{equation}\label{Eq6}
_sA_{lm}=\gamma |c_I|+O(1)\  ,
\end{equation}
where the function $\gamma$ depends on the spin-parameter $s$, the
azimuthal harmonic index $m$, and the spheroidal harmonic index $l$
[see Eq. (\ref{Eq21}) below]. The main aim of the present paper is
to provide a (simple) {\it analytical} proof for the prolate formula
(\ref{Eq6}) in the {\it general} spin case.

\section{A coordinate transformation}

It proves useful to introduce the coordinate $x$ defined by
\cite{Yang,Hodo}
\begin{equation}\label{Eq7}
x\equiv\ln\Big(\tan\Big({{\theta}\over{2}}\Big)\Big)\  ,
\end{equation}
in terms of which the angular equation (\ref{Eq4}) becomes a
Schr\"odinger-like wave equation of the form \cite{NoteSch}
\begin{equation}\label{Eq8}
{{d^2S}\over{dx^2}}-US=0\  ,
\end{equation}
where
\begin{equation}\label{Eq9}
U(\theta)=(m+s\cos\theta)^2-\sin^2\theta(c^2\cos^2\theta-2cs\cos\theta+s+A)\
.
\end{equation}
Note that the interval $\theta\in [0,\pi]$ maps into
$x\in[-\infty,\infty]$. The Schr\"odinger-type angular equation
(\ref{Eq8}) is now in a form that is amenable to a standard WKB
analysis.

\section{The spin-$0$ (scalar) case.}

We shall first consider the spin-0 (scalar) case. In this case the
effective potential $U(\theta)$ is in the form of a symmetric
(invariant under the transformation $\theta\to\pi-\theta$) potential
well: in the $c_I\to\infty$ limit it has a local minimum at
\begin{equation}\label{Eq10}
{_0\theta_{\text{min}}}={\pi\over 2}\ \ \ \text{with} \ \ \
U({_0\theta_{\text{min}}})=-A+O(1)\  .
\end{equation}

Regions where $U(\theta)<0$ are characterized by an oscillatory
behavior of the wave-function $S$ (the `classically allowed
regions'), while regions with $U(\theta)>0$ (the `classically
forbidden regions') are characterized by an exponential behavior
(evanescent waves). The `classical turning points' are characterized
by $U=0$. There is a pair $\{_0\theta^-,_0\theta^+\}$ of such
turning points (with
${_0\theta^-}<{_0\theta_{\text{min}}}<{_0\theta^+}$), which in the
$c_I\to\infty$ limit are located in the near vicinity of
$_0\theta_{\text{min}}$:
\begin{eqnarray}\label{Eq11}
{_0\theta^{\pm}}={\pi\over 2} \pm
{{\sqrt{A}}\over{|c_I|}}+O(c^{-3/2}_I)\ .
\end{eqnarray}

A standard textbook second-order WKB approximation for the
bound-state `energies' of a Schr\"odinger-like wave equation of the
form (\ref{Eq8}) yields the well-known quantization condition
\cite{WKB1,WKB2,WKB3,Iyer,Notehigh}
\begin{equation}\label{Eq12}
\int_{x^{-}}^{x^{+}}dx\sqrt{-U(x)}=(N+{1\over 2})\pi\ \ \ ; \ \ \
N=0,1,2,...\  ,
\end{equation}
where $x^{\pm}$ are the turning points [with $U(x^{\pm})=0$] of the
potential well, and $N$ is a non-negative integer.

Using the relation $dx/d\theta=1/\sin\theta$, one can write the WKB
condition (\ref{Eq12}) in the form
\begin{equation}\label{Eq13}
\int_{{_0\theta^{-}}}^{{_0\theta^{+}}}
d\theta{{\sqrt{-U(\theta)}}\over{\sin\theta}}=(N+{1\over
2})\pi\ \ \ ; \ \ \ N=0,1,2,...\  .
\end{equation}
The WKB quantization condition (\ref{Eq13}) determines the
eigenvalues $\{A\}$ of the associated spin-weighted spheroidal
harmonics in the large-$c_I$ limit. The relation so obtained between
the eigenvalues and the parameters $c,m,s$ and $N$ is rather complex
and involves elliptic integrals. However, given the fact that in the
$c_I\to\infty$ limit the turning points $_0\theta^{\pm}$ lie in the
vicinity of $\theta={\pi\over 2}$ [see Eq. (\ref{Eq11})], one can
approximate the integral in (\ref{Eq11}) by \cite{Noteapp}
\begin{eqnarray}\label{Eq14}
\int_{{_0\theta^{-}}}^{{_0\theta^{+}}}d\theta
\sqrt{-c^2_I(\theta-{\pi\over 2})^2+A}=(N+{1\over 2})\pi\ \ \ ; \ \
\ N=0,1,2,...\  .
\end{eqnarray}
Evaluating the integral in (\ref{Eq14}) is straightforward, and one
finds
\begin{equation}\label{Eq15}
A(N)=(2N+1)|c_I|+O(1)\ \ \ ; \ \ \ N=0,1,2,...\
\end{equation}
for the quantized spectrum. This completes our proof for the prolate
asymptotic spectrum in the scalar ($s=0$) case \cite{NoteCas}.

\section{The general spin case.}

We shall now consider the general spin case. In this case the
effective potential $U(\theta)$ is complex-valued. Its minimum is
located at $_s\theta_{\text{min}}={\pi\over
2}+{{is}\over{c_I}}+O(c^{-2}_I)$, while the two turning points are
located at
\begin{eqnarray}\label{Eq16}
_s\theta^{\pm}={\pi\over 2} \pm
{{\sqrt{A}}\over{|c_I|}}+{{is}\over{c_I}}+O(c^{-3/2}_I)\ .
\end{eqnarray}

A natural generalization of the WKB analysis to the case of
complex-valued potentials is provided in \cite{Miller}: the WKB
quantization rule is given by the standard relation
\begin{equation}\label{Eq17}
\int_{_s\theta^{-}}^{_s\theta^{+}}d\theta{{\sqrt{-U(\theta)}}\over{\sin\theta}}=(N+{1\over
2})\pi\ \ \ ; \ \ \ N=0,1,2,...\  ,
\end{equation}
which can be approximated near $\theta={\pi/2}$ by \cite{Noteapp}
\begin{eqnarray}\label{Eq18}
\int_{_s\theta^{-}}^{_s\theta^{+}}d\theta
\sqrt{-c^2_I(\theta-{\pi\over 2})^2-2isc_I(\theta-{\pi\over
2})+A}=(N+{1\over 2})\pi\ \ \ ; \ \ \ N=0,1,2,...\  .
\end{eqnarray}
As emphasized in \cite{Miller}, the integration path between the two
complex turning points $\{_s\theta^-,_s\theta^+\}$ should be chosen
such that
\begin{equation}\label{Eq19}
\Im\{{\sqrt{-U(\theta)}}\}=0
\end{equation}
along the integration contour \cite{Miller}. In the $c_I\to\infty$
limit this requirement is easily fulfilled by a straight line
(parallel to the real $\theta$-axis) which connects the two turning
points (\ref{Eq16}). Substituting $\theta=\phi-{{is}\over{c_I}}$
(where $\phi\in\mathbb{R}$ runs from ${_0\theta^-}$ to
${_0\theta^+}$) into (\ref{Eq18}) and neglecting terms of order
$O(1)$, one finds that along the path (\ref{Eq19}) the integral
(\ref{Eq16}) can be written as
\begin{eqnarray}\label{Eq20}
\int_{{_0\theta^{-}}}^{{_0\theta^{+}}}d\phi
\sqrt{-c^2_I(\phi-{\pi\over 2})^2+A}=(N+{1\over 2})\pi\ \ \ ; \ \ \
N=0,1,2,...\  .
\end{eqnarray}
This yields
\begin{equation}\label{Eq21}
A(N)=(2N+1)|c_I|+O(1)\ \ \ ; \ \ \ N=0,1,2,...\
\end{equation}
for the quantized spectrum. This completes our proof for the prolate
asymptotic spectrum in the general spin case \cite{NoteCas}. It is
worth emphasizing that the {\it analytical} formula (\ref{Eq21}) for
the prolate asymptotic spectrum agrees with the {\it numerical}
results presented in \cite{BCC,BerCarY}.

%agree with the analytical results of \cite{Flam,Meix} for the
%specific case of spin-$0$ (scalar) fields.

\bigskip
\noindent
{\bf ACKNOWLEDGMENTS}
\bigskip

This research is supported by the Carmel Science Foundation. I thank
Uri Keshet, Oded Hod, Yael Oren, Arbel M. Ongo and Ayelet B. Lata
for helpful discussions.

\end{document}